The sum of two numbers is 18. Twice the smaller number decreased by three equals the larger number. What are the two numbers?
Answer by addingup(3677)

x+y = 18 subtract x on both sides y = 18-x
2x-3 = y substitute for y
2x-3 = 18-x
3x = 21
x = 7
:
John